### v1

# Delayed-feedback oscillators replicate the dynamics of multiplex networks: wavefront propagation and stochastic resonance

Abstract
The widespread development and use of neural networks have significantly enriched a wide range of computer algorithms and promise higher speed at lower cost. However, the imitation of neural networks by means of modern computing substrates is highly inefficient, whereas physical realization of large scale networks remains challenging. Fortunately, delayed-feedback oscillators, being much easier to realize experimentally, represent promising candidates for the empirical implementation of neural networks and next generation computing architectures. In the current research, we demonstrate that coupled bistable delayed-feedback oscillators emulate a multilayer network, where one single-layer network is connected to another single-layer network through coupling between replica nodes, i.e. the multiplex network. We show that all the aspects of the multiplexing impact on wavefront propagation and stochastic resonance identified in multilayer networks of bistable oscillators are entirely reproduced in the dynamics of time-delay oscillators. In particular, varying the coupling strength allows suppressing and enhancing the effect of stochastic resonance, as well as controlling the speed and direction of both deterministic and stochastic wavefront propagation. All the considered effects are studied in numerical simulations and confirmed in physical experiments, showing an excellent correspondence and disclosing thereby the robustness of the observed phenomena.
bistability, noise, wavefront propagation, stochastic resonance, time delay, multilayer network, multiplexing
pacs:
02.30.Ks, 05.10.-a, 05.40.-a, 84.30.-r
The presence of sufficiently long delay in single oscillators can lead to the appearance of the complex high-dimensional dynamics and provide for the observation of oscillatory regimes originally revealed in spatially-extended systems or ensembles of coupled oscillators Yanchuk and Giacomelli (2017). In such a case, spatio-temporal phenomena are tracked down in the purely temporal dynamics of delayed-feedback oscillators by applying a spatio-temporal representation, which considers the delay interval $[0:\tau]$ in analogy with the spatial coordinate Arecchi et al. (1992); Giacomelli and Politi (1996). Using this approach, delay dynamics has been found to sustain stable chimera states Larger et al. (2013, 2015); Semenov et al. (2016); Brunner et al. (2018a), dissipative solitons Garbin et al. (2015); Marconi et al. (2015); Romeira et al. (2016); Brunner et al. (2018a); Semenov and Maistrenko (2018); Yanchuk et al. (2019); Semenov et al. (2023a), travelling waves Klinshov et al. (2017), coarsening Giacomelli et al. (2012); Semenov and Maistrenko (2018) and nucleation Giacomelli et al. (2013). The capability of delay systems to reproduce the high-dimensional phenomena has revolutionized the hardware implementation of recurrent neural networks. In particular, oscillators forced by delayed feedback can efficiently be exploited for reservoir computing as a single-node reservoir Appeltant et al. (2011); Martinenghi et al. (2012); Larger et al. (2017); Brunner et al. (2018b); Duport et al. (2012); Hülser et al. (2022); Köster et al. (2022). Moreover, time-delay reservoirs extended to deep networks can be implemented physically based on the dynamics of coupled nonlinear delayed-feedback oscillators Penkovsky et al. (2019) or a single system with multiple time-delayed feedback loops Stelzer et al. (2021). In addition, one can apply singleton time-delay systems as a spin network (or spin glass) for solving problems with non-deterministic polynomial-time hardness (so-called NP-hard problems). Particularly, Ising machines based on delayed-feedback bistable systems have broad prospects for acceleration of the optimization problem computations Böhm et al. (2019).
Figure 1: (a) Schematic illustration of the dynamics of coupled time delay systems; (b) Circuit diagram of the experimental setup (Eqs. (2)).
In contrast to the single delayed-feedback oscillators, coupled long-delay systems remain unexplored in the context of analogy to the ensembles and networks (except of Ref. Penkovsky et al. (2019) where the deep reservoir scheme comprises hierarchically coupled nonlinear delay oscillators). Meanwhile, the high-dimensional dynamics of coupled oscillators with long delay requires thorough analysis of a number of issues on the role of such coupling as well as on its interpretation in the context of emulated network topology. An intriguing question arises. How accurately does the dynamics of coupled delay oscillators replicate the collective behaviour in complex networks of coupled oscillators? In the current letter, we address this question and demonstrate that coupled delayed-feedback oscillators qualitatively reproduce the dynamics of multilayer networks where the interaction between the layers is realized through the coupling between replica nodes, i.e. multiplex networks Boccaletti et al. (2014); Kivelä et al. (2014). Indeed, in the absence of coupling, delay systems are themselves equivalent to a ring of coupled oscillators with unidirectional coupling. In the presence of diffusive coupling, the oscillators affect each other at any time through the instantaneous values of dynamical variables. The described structure schematically illustrated in Fig. 1 (a) resembles a two-layer multiplex network.
To prove the statement about the similarity to multiplex networks, we show that the intrinsic peculiarities of such networks can be recognized in the dynamics of coupled time-delay oscillators. The distinguishable character of the multiplexing impact on the network dynamics has been repeatedly reported and is responsible for various nontrivial phenomena and control schemes: intra-layer synchronization Gambuzza et al. (2015), control of deviations from synchronization in presence of large perturbations Olmi et al. (2024), reaching the inter-layer synchronization at lower inter-layer coupling strength Eser et al. (2021) and multiplexing-noise-induced synchronization Rybalova et al. (2022), control of chimera Zakharova (2020) and solitary Schülen et al. (2021) states and transitions between them Mikhaylenko et al. (2019); Rybalova et al. (2021), control of coherence Masoliver et al. (2021); Semenova and Zakharova (2018) and stochastic Semenov and Zakharova (2022) resonances and the wavefront propagation Semenov et al. (2023b), topological-asymmetry-induced stabilization of chaotic dynamics (so-called asymmetry-induced order) Medeiros et al. (2021). Of particular interest are the phenomena occurring in a broad spectrum of bistable dynamical systems, namely the wavefront propagation Semenov et al. (2023b) and stochastic resonance Semenov and Zakharova (2022). Both of them are observed and can be efficiently controlled in multiplex networks.
The first effect, wavefront propagation, results from the presence of the asymmetry and can be controlled in a wide variety of dynamical systems including multiplex networks Semenov et al. (2023b) by varying the system parameters or by applying multiplicative noise Engel (1985); Löber et al. (2014); Löcher et al. (1998). The second effect, stochastic resonance Gammaitoni et al. (1998); Neiman et al. (1998); Anishchenko et al. (1999), is manifested through the noise-induced regularity of the stochastic dynamical system response to an input signal and can be enhanced or suppressed by varying the multiplexing strength Semenov and Zakharova (2022). Both phenomena are explored on the example of a system of two coupled bistable delayed-feedback oscillators:
$\begin{array}[]{l}\dfrac{dx}{dt}=-x(x-a_{x})(x+b_{x})+\gamma_{x}(x_{\tau}-x)\\ +\gamma(y-x)+F_{x}(t),\\ \dfrac{dy}{dt}=-y(y-a_{y})(y+b_{y})+\gamma_{y}(y_{\tau}-y)\\ +\gamma(x-y)+F_{y}(t),\\ \end{array}$ (1)
where $\tau$ is the delay time, $x_{\tau}$ and $y_{\tau}$ are the values of variables $x$ and $y$, respectively, at the time moment $(t-\tau)$ with $\tau=1000$. The parameters $\gamma_{x,y}=0.2$ characterize the delayed-feedback strength considered as an analog of the intra-layer coupling strength in the multiplex networks. The coupling strength $\gamma$ represents the strength of the inter-layer interaction in the multiplex network also called the strength of multiplexing (schematically illustrated in Fig. 1 (a)). The external influences are given by $F_{x,y}$. The parameters $a_{x,y}>0$ and $b_{x,y}>0$ define whether the oscillators nonlinearity in the bistable regime is symmetric ($a_{x}=b_{x}$ and $a_{y}=b_{y}$) or asymmetric ($a_{x}\neq b_{x}$ and $a_{y}\neq b_{y}$).
Our investigations are performed by means of numerical simulations and electronic experiments. In more detail, we integrate stochastic differential equations (1) numerically using the Heun method Mannella (2002). For physical experiments, we have developed an experimental prototype being an electronic model of the system (1) implemented by principles of analog modelling Luchinsky et al. (1998). The circuit diagram of the setup is shown in Fig. 1 (b). The detailed description of the circuit is presented in the supplemental material sup . The experimental setup is described by the equations
$\begin{array}[]{l}RC\dfrac{dx}{dt}=-x(x-a_{x})(x+b_{x})+\gamma_{x}(x_{\tau}-x)\\ +\gamma(y-x)+F_{x}(t),\\ RC\dfrac{dy}{dt}=-y(y-a_{y})(y+b_{y})+\gamma_{y}(y_{\tau}-y)\\ +\gamma(x-y)+F_{y}(t),\\ \end{array}$ (2)
where $R=10k\Omega$, $C=10$ nF, $\gamma_{x}=\gamma_{y}=0.2$, $\gamma=0.1V_{*}R/R_{*}$, $\tau=27$ ms. Numerically and experimentally obtained time realizations $x(t)$ and $y(t)$ are mapped onto space-time ($\sigma,n$) by introducing $t=n\eta+\sigma$ with an integer time variable $n$, and a pseudo-space variable $\sigma\in[0,\eta]$, where $\eta=\tau+\varepsilon$ with a quantity $\varepsilon$, which is small compared to $\tau$ and results from a finite internal response time of the system (see the estimation of the resulting $\eta$ in the supplemental material sup ).
Figure 2: Stochastic wavefront propagation control in a single delayed-feedback oscillator $y$ (see Eqs. (1) and (2) by applying parametric noise $\xi_{b}(t)$: $b_{y}=0.45+\xi_{b}(t)$): (a) Dependence of the wavefront propagation speed on the stochastic influence’s variance registered in numerical modelling and electronic experiments. Solid lines are obtained from a fit to the data using the least squares method: $v_{\text{exp}}=1.455-5.078$Var($\xi_{b}(t)$) (red solid line) and $v_{\text{sim}}=0.369-0.756$Var($\xi_{b}(t)$) (blue solid line). Space-time diagrams in panels (b) and (c) illustrate the stochastic dynamics in points 1-3 in numerical and experimental dependencies on panel (a). The upper insets in panels (b) and (c) show the oscillator’s state at $n=10^{3}$. Other parameters: $a_{y}=0.5$, $\gamma_{y}=0.2$ ,$\gamma=0$, $\tau=1000$ (simulations) and $\tau=0.027$ ms (experiments), $F_{x,y}(t)\equiv 0$.
Figure 3: Control of deterministic and stochastic wavefront propagation by varying the coupling $\gamma$. Panel (a) illustrates the interaction of symmetric and asymmetric oscillators ($a_{x}=a_{y}=0.5$, $b_{x}=0.45$, $b_{y}=0.5$), whereas panel (b) corresponds to interaction of two asymmetric oscillators, one of which contains a noise source to stabilize the front propagation as in panels 2 of Fig. 2 (b),(c) ($a_{x}=a_{y}=0.5$, $b_{x}=0.45$, $b_{y}=0.45+\xi_{b}(t)$). Space-time diagrams in panels (c) and (d) illustrate dynamics in points 1-4 in numerical and experimental dependencies on panel (b). Parameters $\gamma_{x,y}$ and $\tau$ are the same as in Fig. 2. External driving is absent, $F_{x,y}(t)\equiv 0$.
Numerical simulations and experiments with analog circuit have shown that the studied delayed-feedback oscillators exhibit propagating fronts in the presence of asymmetry. Thus, the simplest way to control the front propagation speed is varying parameters $a_{x,y}$ and $b_{x,y}$. The second approach, stochastic control by applying multiplicative noise, is illustrated in Fig. 2 on an example of single oscillator $y$ in the presence of asymmetry and parametric Gaussian white noise sup , and in the absence of external influence: $a_{y}=0.5$, $b_{y}=0.45+\xi_{b}(t)$, $F_{y}(t)\equiv 0$. To quantitatively describe the noise-controlled wavefront propagation, the wavefront propagation speed is introduced as the expansion speed of the state $y(t)=a_{y}$ (the red domain in Fig. 2 (b),(c)) sup . The dependence of the mean front propagation speed on the noise variance obtained in numerical simulations, $<v_{\text{sim}}(\text{Var}(\xi_{b}(t)))>$, and electronic experiments, $<v_{\text{exp}}(\text{Var}(\xi_{b}(t)))>$, clearly indicates the option for noise-sustained stabilization and reversal of the wavefront propagation [Fig. 2 (a)]. In the presence of weak noise the dynamics is almost the same as in the deterministic case where one observes expansion of the domain corresponding to the state $y(t)=a_{y}=0.5$ [Fig. 2 (b),(c), panels 1]. For an appropriate noise intensity, such propagation can be stopped (stabilized) by parametric noise [Fig. 2 (b),(c), panels 2] exactly as in spatially-extended systems Engel (1985) and ensembles Semenov et al. (2023b) and then made inverse [Fig. 2 (b),(c), panels 3] for further increasing noise intensity. The opposite effect is achieved if instead modulating parameter $b_{y}$, the parametric noise is introduced for the parameter $a_{y}$ in the form: $a_{y}=0.5+\xi_{a}(t)$. In such a case, increasing noise intensity speeds up the front propagation and the state $y=a_{y}$ invades the available space faster than in the deterministic case.
The results in Fig. 2 (a) indicate the linear character of the dependence of wavefront propagation speed on the noise signal variance. To illustrate this fact, we used curve-fitting based on the least squares method (see the solid lines in Fig. 2 (a)). Slight deviations from the linear functional relationship observed in numerical model result from more pronounced fluctuations of the wavefronts as compared to experimentally registered ones (compare space-time plots in panels 2 of Fig. 2 (b),(c)). To minimize the impact of this effect, we use the mean front propagation speed for the stochastic dynamics description.
It is important to note that the stochastic control of wavefront propagation in single time-delayed feedback oscillators in itself is a novel phenomenon. The fact that multiplicative noise has exactly the same impact on wavefront propagation in bistable media, ensembles and time-delay oscillators with long delay emphasizes the similarity between these classes of high-dimensional dynamical systems. The multiplicative-noise-based control scheme can be applied in addition to deterministic approaches based on tuning the system parameters or applying external periodic signals (one of such methods is successfully applied in delayed optical system in Ref. Marino et al. (2014)).
Next, the dynamics of non-identical coupled oscillators for varying coupling strength $\gamma$ is analysed. In the deterministic case, the parameters are chosen such that oscillator $x$ is asymmetric ($a_{x}=0.5$, $b_{x}=0.45$), while the second oscillator, $y$, is symmetric ($a_{y}=b_{y}=0.5$). In such a case, the front propagation speeds in oscillators $x$ and $y$ approach each other and tend to identical non-zero value when the coupling strength is increased [Fig. 3 (a)]. Once the front propagation speeds have become identical, they do not vary for further increasing $\gamma$ in numerical experiments. In contrast, experimental setup demonstrates variable front propagation speed with increasing $\gamma$ after the common regime is achieved. This is due to the factors inevitably present in experimental electronic setups. The experimental setup is sensitive to inaccuracies which influence the system’s symmetry and hence the wavefront propagation speed. All the circuit active elements, such as operation amplifiers and analog multipliers, have non-zero output offsets. For this reason, it is impossible to realize an absolutely symmetric deterministic oscillator characterised by zero wavefront propagation speed. Moreover, the impact of offsets becomes stronger when increasing $\gamma$, especially when the increase of $\gamma$ is realized by decreasing the resistors $R_{*}$ (see Fig. 1).
In the same way, increasing coupling strength reduces the stochastic dynamics to a common regime. The front propagation speed in the interacting oscillators reaches the same value when both oscillators are asymmetric and the wavefront propagation in oscillator $y$ is stabilised by noise as in panels 2 of Fig. 2 (b),(c) ($a_{x}=a_{y}=0.5$ and $b_{x}=0.45$, $b_{y}=0.45+\xi_{b}$). In contrast to the deterministic oscillator interaction, the front propagation speeds continue to vary with increasing $\gamma$ after the common regime is reached [Fig. 3 (b)-(d)]. This provides for stabilization and reversal of the wavefront propagation in interacting oscillators by varying $\gamma$ both in numerical and physical experiments.
Figure 4: Stochastic resonance in a single delayed-feedback oscillator (oscillator $y$ in Eqs. (1) and (2) subject to periodic forcing and additive noise, $F_{y}(t)=A\sin(\omega t)+\xi_{y}(t)$: (a) Signal-to-noise ratio as a function of the variance of noise. Space-time diagrams in panels (b) and (c) illustrate the stochastic dynamics in points 1-3 in numerical and experimental dependencies in panel (a). The upper insets in panels (b) and (c) show the oscillator’s state at $n=10^{3}$. The system’s parameters: $a_{y}=0.5$, $b_{y}=0.495$, $\gamma_{y}=0.2$, $\gamma=0$, $\tau=1000$ (simulations) and $\tau=0.027$ ms (experiments). External forcing parameters: $A=0.06$, $\omega=0.012$ (simulations) and $A=0.075$, $\omega=448.57$ (physical experiments).
Now let us change the focus and consider single oscillator $y$ subject to external periodic forcing and additive Gaussian white noise, $F_{y}(t)=A\sin(\omega t)+\xi_{y}(t)$, to analyse the occurrence of stochastic resonance [Fig. 4]. The oscillator’s nonlinearity is slightly asymmetric $a_{y}=0.5$, $b_{y}=0.495$, which allows to avoid the impact of multistability. The period of the driving signal is close to half a delay time. The amplitude $A$ is smaller than the threshold value corresponding to the transformation of the dynamics into large-amplitude periodic jumps between two states. Under these conditions, the weak-noise-induced transitions between basins of attraction of two stable steady states $y^{*}_{1}=a_{y}$ and $y^{*}_{2}=b_{y}$ are infrequent (see panels 1 in Fig. 4 (b),(c)). Most of the time the system oscillates near the steady state $y^{*}_{1}=a_{y}$. Noise of larger intensity leads to more and more frequent transitions between the basins of attraction. For an appropriate noise intensity, the transitions are the most regular and the corresponding space-time plots clearly indicate the periodicity in the dynamics of the non-autonomous oscillator [Fig. 4 (b),(c), panels 2]. Increasing noise intensity even further leads to the descrease of the regularity of the systems response [panels 3 in Fig. 4 (b),(c)]. To quantitatively describe the observed noise-induced dynamics, we introduce the signal-to-noise ratio (SNR) sup . The consideration of the SNR as a function of the noise variance in numerical and electronic experiments allows to obtain a non-monotonic curve being a signature of stochastic resonance: there exists an appropriate noise intensity corresponding to the maximal SNR [Fig. 4 (a)].
Figure 5: Enhancement (panels (a) and (b)) and suppression (panels (c) and (d)) of stochastic resonance by varying the coupling strength. Panels (a) and (b) illustrate the SNR for oscillations $y(t)$, whereas the SNRs corresponding to $\gamma=0.03$ and $\gamma=0.08$ in panels (c) and (d) are represented by two curves: the SNR characterizing the oscillations $x(t)$ (lower curves) and the oscillations $y(t)$ (the upper curves). External forces are: $F_{x}(t)=F_{y}(t)=A\sin(\omega t)+\xi_{y}(t)$ (panels (a) and (b)), $F_{x}(t)\equiv 0$ and $F_{y}(t)=A\sin(\omega t)+\xi_{y}(t)$ (panels (c) and (d)). The parameter values are the same as in Fig. 4 except of $\gamma$.
If bistable delayed-feedback oscillators are coupled and forced by a common periodic signal and a source of additive Gaussian white noise, the effect of stochastic resonance can be enhanced. As demonstrated in Fig. 5 (a),(b), one achieves higher signal-to-noise ratios in comparison to a single-system dynamics by varying the coupling strength. However, further increasing coupling returns the stochastic resonance manifestation to the initial, single-system form (except of horizontal shift). Note that the impact of coupling has the resonant character: there is an appropriate coupling strength corresponding to the most pronounced stochastic resonance.
In contrast, stochastic resonance can be suppressed by coupling, if either oscillator is not under direct action of noise and the periodic force [Fig. 5 (c),(d)]. Since there is no interaction between oscillators $x$ and $y$ at $\gamma=0$, for this case Fig. 5 (c),(d) displays the dependence of the SNR on the noise variance only for the oscillator $y$. At non-zero $\gamma$, the curves SNR(Var($\xi_{y}(t)$)) are illustrated for both oscillators $x$ and $y$. Interestingly, we observe coupling-induced stochastic resonance in oscillator $x$, which receives periodic action and noise only through the coupling. However, increasing the coupling strength makes the effect of stochastic resonance in both oscillators less and less pronounced.
In conclusion, we find that the dynamics of coupled delayed-feedback oscillators explored in numerical simulations and physical experiments reproduces with high accuracy the effects in multiplex networks of bistable elements. In the context of this similarity, the delayed feedback strength plays role of the coupling within the layer (intra-layer coupling), whereas the coupling between the two delayed-feedback oscillators corresponds to the multiplex links between the layers (inter-layer coupling).
Intriguingly, all the aspects of the multiplexing-based wavefront propagation control reported in Ref. Semenov et al. (2023b) are replicated by a numerical model of two oscillators with time delay. This is manifested through the reduction of the wavefront propagation to a common regime where the front propagation speeds tend to the same value. This value is fixed in the deterministic model, but varies in the stochastic oscillators with increasing the strength of multiplexing. This provides for stabilization and reversal of the front propagation in the interacting oscillators. It is important to note that due to the specific properties of the experimental setup, the common regime where the propagation speed is fixed with increasing coupling strength can not be achieved. In addition, the intrinsic peculiarities of the multiplexing-based control of stochastic resonance established in Ref. Semenov and Zakharova (2022) have been revealed in a pair of coupled delayed-feedback oscillators. In more detail, coupling suppresses stochastic resonance, if the periodic forcing and noise are present in only one of the two oscillators. In contrast, increasing the coupling strength allows to enhance the stochastic resonance, if the periodic forcing and noise are present in all the interacting systems. In such a case, the impact of multiplexing has a resonant character.
The presented results have a wide spectrum of potential applications in the context of the delayed-feedback-oscillator-based physical implementations of machine learning algorithms initially realized by means of multilayer networks Penkovsky et al. (2019); Stelzer et al. (2021), especially the multiplex ones Amoroso et al. (2019). In addition, a distinguishable application where the property of bistability and the symmetry control are of particular importance is the physical implementation of spin-networks as solvers of combinatorial optimization problems Böhm et al. (2019); Mohseni et al. (2022).
This work has been supported by the Russian Science Foundation (project No. 22-72-00038).